# Fortify the Guardian, Not the Treasure: Resilient Adversarial Detectors

### v1

@@ 5 Results
Initially, we conducted an assessment of classifier performance following the generation of adversarial perturbations employing the PGD method. The outcomes of this evaluation are delineated in Table 1, encompassing classifier performance metrics both on clean and perturbed test sets. It is evident that the classifiers exhibit considerable susceptibility to adversarial manipulations, with the crafted attacks achieving near-perfect efficacy in deceiving the classifiers.
After training involving the use of PGD-generated adversarial images, Table 2 shows the performance outcomes of the adversarial detectors, as assessed through the ROC-AUC metric, before the deployment of RADAR.
Table 2: Without RADAR: Performance of detectors on several classifiers and datasets. In this and subsequent tables, boldface marks best performance. Note: for ROC AUC, higher is better.
CIFAR-10 SVHN
Detector $\text{AUC}_{\text{Avg.}}$ $\text{AUC}_{\text{PGD}}$ $\text{AUC}_{\text{OPGD}}$ $\text{AUC}_{\text{SPGD}}$ $\text{AUC}_{\text{PGD}}$ $\text{AUC}_{\text{OPGD}}$ $\text{AUC}_{\text{SPGD}}$
VGG-11 0.33 1.00 0.00 0.00 1.00 0.00 0.00
VGG-13 0.33 1.00 0.00 0.00 1.00 0.00 0.00
VGG-16 0.53 1.00 0.61 0.59 1.00 0.00 0.00
ResNet-18 0.33 1.00 0.00 0.00 1.00 0.00 0.00
ResNet-34 0.33 1.00 0.00 0.00 1.00 0.00 0.00
ResNet-50 0.39 1.00 0.15 0.14 1.00 0.03 0.05
Avg. 0.37 1.00 0.13 0.12 1.00 0.00 0.00
Table 3 presents results regarding the adversarial detectors’ performance, based on the SR@5 metric, prior to using RADAR.
@@
VGG-11 0.97 0.97 0.97 0.98 0.99
VGG-13 0.98 0.97 0.99 0.99 0.99
VGG-16 0.68 0.37 0.40 0.99 0.99
ResNet-18 0.97 0.96 0.97 0.99 0.99
ResNet-34 0.97 0.96 0.97 0.99 0.99
ResNet-50 0.88 0.81 0.83 0.96 0.94
Avg. 0.90 0.84 0.85 0.98 0.98
We then deployed RADAR. The resultant outcomes pertaining to the efficacy of the adversarial detectors, as measured by the ROC-AUC metric, after the integration of the RADAR framework, are delineated in Table 4.
@@
CIFAR-10 SVHN
Detector $\text{AUC}_{\text{Avg.}}$ $\text{AUC}_{\text{PGD}}$ $\text{AUC}_{\text{OPGD}}$ $\text{AUC}_{\text{SPGD}}$ $\text{AUC}_{\text{PGD}}$ $\text{AUC}_{\text{OPGD}}$ $\text{AUC}_{\text{SPGD}}$
VGG-11 0.98 0.99 0.95 0.94 1.00 1.00 1.00
VGG-13 0.99 0.99 0.99 0.99 1.00 1.00 1.00
VGG-16 0.99 1.00 0.99 0.99 1.00 1.00 1.00
ResNet-18 0.98 0.99 0.96 0.96 1.00 0.99 1.00
ResNet-34 0.99 0.99 0.99 1.00 1.00 1.00 1.00
ResNet-50 0.99 0.99 0.99 0.99 1.00 1.00 0.99
Avg. 0.99 0.99 0.98 0.98 1.00 0.99 0.99
Table 5 presents results regarding the adversarial detectors’ performance, based on the SR@5 metric, upon deploying RADAR.
@@
VGG-11 0.02 0.04 0.06 0.00 0.00
VGG-13 0.00 0.00 0.00 0.00 0.00
VGG-16 0.00 0.00 0.00 0.00 0.00
ResNet-18 0.02 0.05 0.05 0.00 0.00
ResNet-34 0.00 0.00 0.00 0.00 0.00
ResNet-50 0.00 0.00 0.00 0.00 0.00
Avg. 0.00 0.01 0.02 0.00 0.00
Table 6 and Table 7 show the generalization performance of RADAR-trained detectors on classifiers they were not trained on.
Table 6: Generalization performance of adversarially trained detectors, trained on CIFAR-10, on different classifiers. A value represents the ROC AUC of the respective detector/classifier pair, for OPGD and SPGD, respectively. For example, the two top-right values represent the ROC AUC for OPGD and SPGD, of a ResNet50 detector model, trained on a ResNet50 classifier model, with its performance then evaluated on attacks involving a VGG-11 classifier model. Clf: Classifier model. Det: Adversarial detector model.
Clf Det VGG-11 VGG-13 VGG-16 ResNet-18 ResNet-34 ResNet-50
VGG-11 0.95, 0.94 0.99, 0.99 0.99, 0.99 0.96, 0.99 0.99, 0.99 0.99, 0.99
VGG-13 0.95, 1.00 0.99, 0.99 0.99, 0.99 0.96, 0.99 0.99, 0.99 0.99, 0.99
VGG-16 0.95, 0.95 0.99, 0.99 0.99, 0.99 0.96, 0.99 0.99, 0.99 0.99, 0.99
ResNet-18 0.95, 0.94 0.99, 0.99 0.99, 0.99 0.96, 0.96 0.99, 0.99 0.99, 0.99
ResNet-34 0.95, 0.94 0.99, 0.99 0.99, 0.99 0.96, 0.99 0.99, 1.00 0.99, 0.99
ResNet-50 0.95, 0.94 0.99, 0.99 0.99, 0.99 0.96, 0.99 0.99, 0.99 1.00, 0.99
Table 7: Generalization performance of adversarially trained detectors, trained on SVHN, on different classifiers. A cell value represents the ROC AUC of the respective detector/classifier pair. Clf: Classifier model. Det: Adversarial detector model.
Clf Det VGG-11 VGG-13 VGG-16 ResNet-18 ResNet-34 ResNet-50
VGG-11 1.00, 1.00 1.00, 1.00 1.00, 1.00 0.99, 0.99 1.00, 1.00 1.00, 0.99
VGG-13 1.00, 1.00 1.00, 1.00 1.00, 1.00 0.99, 0.99 1.00, 1.00 1.00, 0.99
VGG-16 1.00, 1.00 1.00, 1.00 1.00, 1.00 0.99, 0.99 1.00, 1.00 1.00, 0.99
ResNet-18 1.00, 1.00 1.00, 1.00 1.00, 1.00 0.99, 1.00 1.00, 1.00 1.00, 0.99
ResNet-34 1.00, 1.00 1.00, 1.00 1.00, 1.00 0.99, 0.99 1.00, 1.00 1.00, 0.99
ResNet-50 1.00, 1.00 1.00, 1.00 1.00, 1.00 0.99, 0.99 1.00, 1.00 1.00, 0.99
@@
Before the incorporation of adversarial training, the optimization process exhibited a trend of rapid convergence towards zero loss within a few iterations, as can be seen in the top rows of Figure 2, Figure 3, Figure 4, and Figure 5. Once we integrated adversarial training into the detector, we observed a distinct shift in the optimization behavior.
Figure 2: CIFAR-10. Binary cross-entropy loss metrics, from the point of view of an attacker, are herein presented in the context of crafting an adversarial instance from the test set, using VGG-11 (left), VGG-13 (middle), and VGG-16 (right). These plots illustrate the progression of loss over 20 different images of orthogonal projected gradient descent (OPGD), with the main goal being to minimize the loss. Top: Prior to adversarial training, the loss converges to zero after a small number of iterations. Bottom: After adversarial training, the incurred losses are significantly higher by orders of magnitude (note the difference in scales), compared to those observed in their standard counterparts. This shows that the detector is now resilient, i.e., far harder to fool.
Figure 3: CIFAR10. Binary cross-entropy loss metrics, from the point of view of an attacker, are herein presented in the context of crafting an adversarial instance from the test set using ResNet-18 (left), ResNet-34 (middle), and ResNet-50 (right).
Figure 4: SVHN. Binary cross-entropy loss metrics, from the point of view of an attacker, are herein presented in the context of crafting an adversarial instance from the test set using VGG-11 (left), VGG-13 (middle), and VGG-16 (right).
Figure 5: SVHN. Binary cross-entropy loss metrics, from the point of view of an attacker, are herein presented in the context of crafting an adversarial instance from the test set using ResNet-18 (left), ResNet-34 (middle), and ResNet-50 (right).
@@
Upon deploying RADAR the optimization process displayed a tendency to plateau after a small number of iterations, with loss values typically higher by orders of magnitude, as compared to those observed prior to deployment. This plateau phase persisted across diverse experimental settings and datasets, indicating a fundamental change in the optimization landscape, induced by RADAR.
This behavior can be seen as a robustness-enhancing effect, because the detector appears to resist rapid convergence towards trivial solutions, thereby enhancing its generalization capabilities. This showcases the efficacy of our method in bolstering the defenses of detection systems against adversarial threats—without sacrificing clean accuracy.

### v2

@@ 5 Results
Initially, we conducted an assessment of classifier performance following the generation of adversarial perturbations employing the PGD method.
The outcomes of this evaluation, presented in Table 1, highlight classifier performance on both clean and perturbed test sets. The results reveal significant vulnerability to adversarial manipulations. For instance, on the CIFAR-10 dataset, the VGG-11 classifier drops from 92.40% accuracy on clean data to 0.35% on adversarial examples. Other classifiers on CIFAR-10 show similar trends, with clean accuracy values between 92.40% and 94.21%, and adversarial accuracy values below 6%. A similar pattern is observed on the SVHN dataset, where the VGG-11 classifier’s accuracy falls from 93.96% on clean data to 0.00% on adversarial attacks. Other SVHN classifiers also exhibit significant performance drops, with adversarial accuracy values near 0%. On the ImageNet dataset, all classifiers reached 0% accuracy when attacked. These findings illustrate the stark vulnerability of standard classifiers to adversarial perturbations, underscoring the need for more robust defense mechanisms.
After standard adversarial training involving the use of PGD-generated adversarial images, Table 2 shows the performance outcomes of the adversarial detectors, as assessed through the ROC-AUC metric, before the deployment of RADAR. The table summarizes the average ROC-AUC ($\text{AUC}_{\text{Avg.}}$) and individual ROC-AUC scores for PGD ($\text{AUC}_{\text{PGD}}$), OPGD ($\text{AUC}_{\text{OPGD}}$), and SPGD ($\text{AUC}_{\text{SPGD}}$) attacks. For the CIFAR-10 dataset, VGG-16 detector achieved the highest average ROC-AUC score of 0.46, which is slightly worse than tossing a coin. All detectors performed perfectly against PGD attacks, but their performance dropped significantly against OPGD and SPGD, with most detectors showing a score close to 0.00. The SVHN and ImageNet datasets displayed a similar trend. All models achieved perfect detection against PGD attacks, but their effectiveness plummeted for OPGD and SPGD attacks, with only ResNet-50 showing minimal detection capability.
Table 2: Without RADAR: Performance of detectors on several classifiers and datasets. In this and subsequent tables, boldface marks best performance. Note: for ROC-AUC, higher is better.
Dataset $\text{AUC}_{\text{Avg.}}$ CIFAR-10 SVHN ImageNet
$\text{AUC}_{\text{PGD}}$ $\text{AUC}_{\text{OPGD}}$ $\text{AUC}_{\text{SPGD}}$ $\text{AUC}_{\text{PGD}}$ $\text{AUC}_{\text{OPGD}}$ $\text{AUC}_{\text{SPGD}}$ $\text{AUC}_{\text{PGD}}$ $\text{AUC}_{\text{OPGD}}$ $\text{AUC}_{\text{SPGD}}$
VGG-11 0.33 1.00 0.00 0.00 1.00 0.00 0.00 1.00 0.00 0.00
VGG-13 0.33 1.00 0.00 0.00 1.00 0.00 0.00 1.00 0.00 0.00
VGG-16 0.46 1.00 0.61 0.59 1.00 0.00 0.00 1.00 0.00 0.00
ResNet-18 0.33 1.00 0.00 0.00 1.00 0.00 0.00 1.00 0.00 0.00
ResNet-34 0.33 1.00 0.00 0.00 1.00 0.00 0.00 1.00 0.00 0.00
ResNet-50 0.37 1.00 0.15 0.14 1.00 0.03 0.05 1.00 0.00 0.00
Avg. 0.36 1.00 0.13 0.12 1.00 0.00 0.00 1.00 0.00 0.00
[…]
@@
[…]
VGG-11 0.98 0.97 0.97 0.98 0.99 0.99 1.00
VGG-13 0.99 0.97 0.99 0.99 0.99 1.00 1.00
VGG-16 0.78 0.37 0.40 0.99 0.99 0.97 0.99
ResNet-18 0.98 0.96 0.97 0.99 0.99 0.99 1.00
ResNet-34 0.98 0.96 0.97 0.99 0.99 0.99 1.00
ResNet-50 0.93 0.81 0.83 0.96 0.94 0.97 0.99
Avg. 0.90 0.84 0.85 0.98 0.98 0.98 0.99
We then deployed RADAR. The outcomes on the efficacy of adversarial detectors, measured by ROC-AUC and SR@5 metrics after integrating RADAR, are presented in Table 5 and Table 6. Table 4 shows the accuracy percentages of all the classifiers on clean and adversarially perturbed test sets across CIFAR-10, SVHN, and ImageNet datasets, after applying RADAR. Note that the benign accuracy for all classifiers across CIFAR-10, SVHN, and ImageNet datasets remained unchanged. For all datasets, RADAR maintains high accuracy both on PGD and OPGD samples. For example, VGG-11 on CIFAR-10 retains an accuracy of 92.40% across all scenarios, while ResNet-50 on SVHN shows only a slight reduction from 94.33% to 94.32% on OPGD samples. In contrast, the ImageNet dataset reveals more significant accuracy drops under adversarial conditions, particularly for ResNet-18, which falls from 70.24% on benign samples to 61.05% on PGD samples.
Table 4: With RADAR: The accuracy percentage of the original classifiers is assessed both on clean and adversarially perturbed test sets, using both PGD and OPGD, subsequent to the application of RADAR. In instances where the adversarial detector indicates an adversarial sample (adv), the classifier’s prediction is disregarded.
Dataset VGG-11 VGG-13 VGG-16 ResNet-18 ResNet-34 ResNet-50
Ben PGD OPGD Ben PGD OPGD Ben PGD OPGD Ben PGD OPGD Ben PGD OPGD Ben PGD OPGD
CIFAR-10 92.40 92.40 92.37 94.21 94.17 94.08 94.00 93.96 93.91 92.60 92.41 91.84 93.03 93.03 93.03 93.43 93.43 93.14
SVHN 93.96 93.96 93.95 94.85 94.85 94.85 94.95 94.95 94.94 94.88 94.88 94.86 94.84 94.83 94.82 94.33 94.33 94.32
ImageNet 70.08 68.89 68.73 70.44 69.87 69.83 71.96 71.96 71.96 70.24 61.05 68.63 72.24 68.18 72.18 74.96 74.63 74.96
Following RADAR integration, we observed significant improvements in adversarial detection performance across various classifiers and datasets, as shown in Table 5. Notably, on CIFAR-10, models such as VGG-13, VGG-16, ResNet-34, and ResNet-50 achieved high average ROC-AUC scores of 0.99, with ResNet-18 slightly lower at 0.98. For SVHN, all models achieved ROC-AUC scores of 1.00, except for ResNet-18 and ResNet-50, which scored 0.99 under specific attack conditions. Similarly, for the ImageNet dataset, models exhibited robust performance, with most achieving ROC-AUC scores of 1.00. Specifically, ResNet-50 scored consistently high at 1.00 or 0.99 across different attack types. These results underscore RADAR’s ability to fortify adversarial detectors against various attack methods.
The SR@5 metric further confirms the robustness of RADAR-enhanced detectors, as detailed in Table 6. Models like VGG-13, VGG-16, ResNet-34, and ResNet-50 achieved perfect SR@5 scores of 0.00 on all datasets, indicating successful detection of adversarial attacks. VGG-11 and ResNet-18 demonstrated slightly lower performance on CIFAR-10. RADAR significantly improves the robustness of adversarial detectors, as evidenced by high ROC-AUC scores and low SR@5 values across different models and datasets, including the challenging ImageNet dataset. This comprehensive performance across diverse datasets reinforces the efficacy of our proposed approach in enhancing adversarial detector resilience.
@@
Dataset $\text{AUC}_{\text{Avg.}}$ CIFAR-10 SVHN ImageNet
$\text{AUC}_{\text{PGD}}$ $\text{AUC}_{\text{OPGD}}$ $\text{AUC}_{\text{SPGD}}$ $\text{AUC}_{\text{PGD}}$ $\text{AUC}_{\text{OPGD}}$ $\text{AUC}_{\text{SPGD}}$ $\text{AUC}_{\text{PGD}}$ $\text{AUC}_{\text{OPGD}}$ $\text{AUC}_{\text{SPGD}}$
VGG-11 0.98 0.99 0.95 0.94 1.00 1.00 1.00 1.00 1.00 1.00
VGG-13 0.99 0.99 0.99 0.99 1.00 1.00 1.00 1.00 0.99 0.99
VGG-16 0.99 1.00 0.99 0.99 1.00 1.00 1.00 1.00 0.99 0.99
ResNet-18 0.98 0.99 0.96 0.96 1.00 0.99 1.00 0.95 1.00 1.00
ResNet-34 0.99 0.99 0.99 1.00 1.00 1.00 1.00 0.98 1.00 1.00
ResNet-50 0.99 0.99 0.99 0.99 1.00 1.00 0.99 0.99 1.00 1.00
Avg. 0.99 0.99 0.98 0.98 1.00 0.99 0.99 0.98 0.99 0.99
@@
[…]
VGG-11 0.02 0.04 0.06 0.00 0.00 0.00 0.00
VGG-13 0.00 0.00 0.00 0.00 0.00 0.00 0.00
VGG-16 0.00 0.00 0.00 0.00 0.00 0.00 0.00
ResNet-18 0.02 0.05 0.05 0.00 0.00 0.00 0.00
ResNet-34 0.00 0.00 0.00 0.00 0.00 0.00 0.00
ResNet-50 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Avg. 0.00 0.01 0.02 0.00 0.00 0.00 0.00
Figure 2 show the generalization performance of RADAR-trained detectors on classifiers they were not trained on. This table shows the ROC-AUC values of RADAR-trained detectors, when evaluated on the different classifier models. Each cell in the table corresponds to the ROC-AUC value achieved by the detector-classifier pair. For example, the top-left cell in the top-right table indicates the performance of a VGG-11 detector model trained on a VGG-11 classifier model using the ImageNet dataset, and subsequently evaluated on attacks utilizing a ResNet-50 classifier model. The results indicate consistently high generalization across different classifiers, with most detector-classifier pairs achieving ROC-AUC values of 0.99 or higher. This demonstrates the effectiveness of RADAR-trained detectors in maintaining high adversarial detection performance across diverse datasets and classifier architectures. These findings underscore the resilience of adversarial detectors trained with RADAR, showcasing their ability to maintain robust detection capabilities even when confronted with unseen classifiers. The high ROC-AUC values indicate that our approach effectively fortifies the detectors themselves, rather than relying solely on robust classifier training, thereby enhancing the overall security and reliability of the adversarial detection system.
Figure 2: Generalization performance of adversarially trained detectors, trained on CIFAR-10, SVHN and ImageNet. Each adversarial detector was trained using each corresponding classifier, e.g. ResNet-50 adversarial detector was trained using ResNet-50 image classifier. This table shows the generalization of each detector to other classifiers, which it didn’t train with. A value represents the ROC-AUC of the respective detector/classifier pair, for OPGD (top row) and SPGD (bottom row) with $\epsilon=\frac{16}{255}$.
@@
Impact of adversarial training on optimization dynamics. Before the incorporation of adversarial training, the optimization process exhibited a trend of rapid convergence towards zero loss within a few iterations, as can be seen in the top rows of Figure 3, Figure 4 and Figure 5. Once we integrated adversarial training into the detector, we observed a distinct shift in the optimization behavior. Upon deploying RADAR, the optimization process displayed a tendency to plateau after a small number of iterations, with loss values typically higher by orders of magnitude, as compared to those observed prior to deployment. This plateau phase persisted across diverse experimental settings and datasets, indicating a fundamental change in the optimization landscape, induced by RADAR. This behavior can be seen as a robustness-enhancing effect, because the detector appears to resist rapid convergence towards trivial solutions, thereby enhancing its generalization capabilities. This showcases the efficacy of our method in bolstering the defenses of detection systems against adversarial threats—without sacrificing clean accuracy.
Figure 3: CIFAR-10. Binary cross-entropy loss metrics, from the point of view of an attacker, are herein presented in the context of crafting an adversarial instance from the test set. These plots illustrate the progression of loss over 20 different images of orthogonal projected gradient descent (OPGD), with the main goal being to minimize the loss. Top: Prior to adversarial training, the loss converges to zero after a small number of iterations. Bottom: After adversarial training, the incurred losses are significantly higher by orders of magnitude (note the difference in scales), compared to those observed in their standard counterparts. This shows that the detector is now resilient, i.e., far harder to fool.
Figure 4: SVHN. Binary cross-entropy loss metrics, from the point of view of an attacker, are herein presented in the context of crafting an adversarial instance from the test set.
Figure 5: ImageNet. Binary cross-entropy loss metrics, from the point of view of an attacker, are herein presented in the context of crafting an adversarial instance from the test set using OPGD.
Robustness analysis with various epsilon values. The results of our experiments, depicted in Figure 6 and Figure 7, provide a comprehensive evaluation of the robustness of the adversarial detectors against OPGD and SPGD attacks with different $\epsilon$ values.
For the CIFAR-10, SVHN, and ImageNet datasets, increasing $\epsilon$ values generally results in increasing ROC-AUC scores and decreasing SR@5 rates, using both OPGD and SPGD, indicating improved detection capabilities and decreased success of adversarial attacks.
Specifically, VGG-11 shows a significant drop in AUC and a corresponding rise in SR@5 as $\epsilon$ decreases, reflecting its susceptibility to lower magnitude perturbations. VGG-13 and VGG-16 exhibit similar patterns, though VGG-16 demonstrates slightly better robustness at lower $\epsilon$ values. This trend is consistent across both OPGD and SPGD attack evaluations.
The ResNet models, particularly ResNet50, demonstrate superior resilience performance. Across various $\epsilon$ values, ResNet50 consistently exhibits higher ROC-AUC scores and lower SR@5 rates compared to other models, signifying its robust ability to detect subtle adversarial examples. Notably, our method maintains consistently low SR@5 values across all datasets, underscoring its resilience. Specifically, experiments on the SVHN dataset consistently achieve an SR@5 of 0%. For CIFAR-10 and ImageNet, SR@5 ranges between 0% to 17% with OPGD, and between 0% to 15% with SPGD, affirming the efficacy of our approach in enhancing adversarial-detection capability. These results highlight the advantage of applying adversarial training directly to adversarial detectors rather than focusing solely on classifiers themselves.
Figure 6: AUC and SR@5 scores across different epsilon values for CIFAR-10, SVHN, and ImageNet datasets using OPGD. The performance of the adversarial detectors is illustrated, highlighting how AUC and SR@5 varies across different perturbation magnitudes.
Figure 7: AUC and SR@5 scores across different epsilon values for CIFAR-10, SVHN, and ImageNet datasets using SPGD.
Comparison to other detectors against adaptive attacks. The table presents the robust accuracy ($\text{Acc}_{\text{robust}}$) of various defenses against OPGD and SPGD adaptive attacks at two perturbation levels, $\epsilon=0.01$ and $\epsilon=8/255$, and at two false positive rates, 5% (FP@5) and 50% (FP@50). The defenses evaluated are ContraNet (Yang et al., 2022b), Trapdoor (Shan et al., 2020), DLA (Sperl et al., 2020), and SID (Tian et al., 2021). RADAR consistently achieves the highest robust accuracy across both attack types and perturbation levels, maintaining nearly 100% accuracy in all scenarios. ContraNet also shows high performance, particularly at lower perturbation levels, but its accuracy decreases at higher perturbation levels. Trapdoor, DLA, and SID exhibit varying degrees of robustness, with significant decreases in accuracy at higher perturbation levels. For instance, Trapdoor’s accuracy drops to almost 0% under both attack types at $\epsilon=8/255$.
Table 7: Robust accuracy of different defenses under OPGD and SPGD attacks. The table presents the robust accuracy ($\text{Acc}_{\text{robust}}$) of various defenses when subjected to the adaptive attacks, OPGD and SPGD. The accuracy is evaluated at two perturbation levels, $\epsilon=0.01$ and $\epsilon=8/255$, and is reported for two false positive rates (FP), 5% (FP@5) and 50% (FP@50).
Attack Defense $\epsilon=0.01$ $\epsilon=8/255$
$\text{Acc}_{\text{robust}}FP@5$ $\text{Acc}_{\text{robust}}FP@50$ $\text{Acc}_{\text{robust}}FP@5$ $\text{Acc}_{\text{robust}}FP@50$
OPGD RADAR 99.2% 99.3% 99.8% 99.9%
ContraNet (Yang et al., 2022b) 93.7% 99.2% 89.8% 94.7%
Trapdoor (Shan et al., 2020) 0.0% 7.0% 0.0% 8.0%
DLA (Sperl et al., 2020) 62.6% 83.7% 0.0% 28.2%
SID (Tian et al., 2021) 6.9% 23.4% 0.0% 1.6%
SPGD RADAR 99.2% 99.1% 99.8% 99.9%
ContraNet (Yang et al., 2022b) 93.7% 99.3% 89.7% 95.1%
Trapdoor (Shan et al., 2020) 0.2% 49.5% 0.4% 37.2%
DLA (Sperl et al., 2020) 17.0% 55.9% 0.0% 13.5%
SID (Tian et al., 2021) 8.9% 50.9% 0.0% 11.4%
5.1 Ablation Studies
In this section, we conduct comprehensive ablation studies to evaluate the impact of critical hyperparameters in our proposed loss function on the performance of our adversarial detectors on the validation set. Specifically, we focus on four key parameters: number of steps, step size ($\alpha$), batch size, and learning rate. Each of these parameters plays a significant role in the training process and potentially influences the robustness and effectiveness of our adversarial detectors. To evaluate the effectiveness and trade-offs associated with these parameters, we conducted a series of experiments using the ResNet-50 architecture on the CIFAR-10 dataset. The results are delineated in Figure 8.
Figure 8: Ablation studies using VGG-11 with different number of steps, $\alpha$, batch size, and learnig rate values, from left to right.
The results reveal that the number of steps significantly affects the model’s performance. The detector’s performance reaches its peak at 100 steps. This suggests that more iterations in the adversarial training process lead to better optimization and enhanced robustness of the adversarial detector. The choice of $\alpha$ also plays a critical role. Our findings show that a smaller $\alpha$ (0.05) yields the best performance, while increasing $\alpha$ increases effectiveness. This indicates that bigger perturbations during training are more effective in strengthening the detector’s resilience to adversarial attacks. But we want to detect attacks that are unrecognizable, thus we chose to use $\alpha=0.03$.
Regarding batch size, the performance remains relatively stable across smaller batch sizes but shows a notable decline at the largest batch size tested (256).
Learning rate is another crucial factor. Initially, a learning rate of $1\times 10^{-1}$ and $1\times 10^{-2}$ performs worse, but as the learning rate decreases, there is an improvement in performance. However, a learning rate that is too small ($1\times 10^{-5}$) results in under-performance, highlighting the need for a balanced learning rate to ensure effective training.
Based on this ablation study, we selected the hyperparameters that yielded the best performance: a batch size of 32, a learning rate of $1\times 10^{-2}$, 100 steps, and an $\alpha$ value of 0.03. These settings were found to optimize the robustness of our adversarial detector, providing a strong defense against adversarial attacks.